Find the remainder when f(x) is divided by (x - k)

f(x) = 4x3 - 6x2 + 3x + 1; k= -2

-3

-61

26

-49

show work please and explanation
1
Expert's answer
2014-05-26T05:56:43-0400
f(x)/(x+2) = (4x3 - 6x2 + 3x + 1)/(x+2)
4x3 - 6x2 + 3x + 1|x+2
4x3+8x2 |4x^2-14x+31
-14x2+3x
-14x2-28x
31x+1
31x+62
-61
And we get
f(x)/(x+2) = (4x3 - 6x2 + 3x + 1)/(x+2) = (4x2-14x+31)(x+2) - 61
remainder is -61